# Metrics on completely positive maps via noncommutative geometry

### v1

@@ 2.2 Lipschitz seminorms and spectral triples
$\operatorname{mk}_{L}(\phi,\psi):=\sup\{|\phi(a)-\psi(a)|:a\in A,L(a)\leq 1\}$ (2.1)
When $L$ is a Lipschitz seminorm on $A$, there has been a lot of interest in deciding when $\operatorname{mk}_{L}$ metrizes the weak topology on $S(A)$, that is, when the induced topology from this extended metric coincides with the weak-topology. This is a central question when studying compact quantum metric spaces, but we will never explicitly need to consider this question in this article.
@@
A unital spectral triple is even when the Hilbert space $H$ is equipped with a $\mathbb{Z}/2\mathbb{Z}$-grading operator $\gamma\colon H\to H$ such that $\pi(a)\colon H\to H$ is an even operator for all $a\in\mathcal{A}$ and $D\colon{\rm Dom}(D)\to H$ is odd. Otherwise we say the unital spectral triple is odd.
@@
$L_{\partial}(a)=\begin{cases}\|[\partial,a]\|&\text{if $a\in\mathcal{A}$}\\ \infty&\text{if $a\in A\setminus\mathcal{A}$},\end{cases}$ (2.2)
@@
Constructing a Lipschitz seminorm from a spectral triple highlights an important subtlety: We set the domain ${\rm Dom}(L)$ to be equal to $\mathcal{A}$, even if it is possible that the quantity $\|[\partial,a]\|<\infty$ for some $a\in A\setminus\mathcal{A}$. Despite this, we still set $L_{\partial}(a)=\infty$ in this case. All this is to say that the domain of $L$ need not agree with the maximal domain for the derivation $a\mapsto[\partial,a]$.
@@
$\begin{split}(L_{A}\otimes 1)(z)&:=\sup\{L((\operatorname{id}_{A}\otimes\psi)(z))\colon\psi\in S(B)\}\\ (1\otimes L_{B})(z)&:=\sup\{L((\phi\otimes\operatorname{id}_{B})(z))\colon\phi\in S(A)\}\\ L_{A\otimes B}&:=L_{A}\otimes 1+1\otimes L_{B}\end{split}$ (2.3)
where the domain for all three of these seminorms is set to be ${\rm Dom}(L_{A})\odot{\rm Dom}(L_{B})$ which is dense in $A\otimes B$. Assume now that $L_{A}$ and $L_{B}$ are Lipschitz seminorms. Using the fact that $L_{A}$ and $L_{B}$ are both *-invariant, it is not difficult to verify that the three seminorms just defined are *-invariant. Moreover, we easily verify that $\mathbb{C}\cdot(1_{A}\otimes 1_{B})$ is contained in the kernel of all three seminorms. Thus they are all Lipschitz seminorms. Note however that $\ker(L_{A}\otimes 1)$ is generally larger than $\mathbb{C}\cdot(1_{A}\otimes 1_{B})$, as it at the very least contains $\mathbb{C}\cdot(1_{A}\otimes{\rm Dom}(L_{B}))$. The analogous statement holds for $1\otimes L_{B}$. Moreover, $L_{A\otimes B}(x)=L_{B\otimes A}(\Sigma(x))$ for all $x\in A\otimes B$.
@@
3 Discussions on inducing metrics on unital completely positive maps
Throughout this section, $A$ and $B$ will be unital $C^{*}$-algebras, and we will be considering extended metrics on $\operatorname{UCP}(A,B)$, the set of unital completely positive maps from $A$ to $B$. In [GLN05] and [BV25], several desirable properties of such metrics are discussed, as well as some examples of metrics in the case where $A$ and $B$ are matrix algebras. Note that in [GLN05, BV25] they discuss quantum channels between $A$ and $B$ instead of unital completely positive maps. These classes of maps are dual to each other, and for the sake of clarity, we will simply translate the discussions in [GLN05] and [BV25] to how it would fit into our current setting without further mentions.
@@ 3 Discussions on inducing metrics on unital completely positive maps
The first property an extended metric on $\operatorname{UCP}(A,B)$ should satisfy is stability. In both [GLN05] and [BV25], a metric $d$ on $\operatorname{UCP}(A,B)$ is said to be stable if
$d(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)=d(F,G)$ (3.1)
@@
holds for every $n\in\mathbb{N}$ and $F,G\in\operatorname{UCP}(A,B)$. From a quantum information theoretic perspective, this property is supposed to reflect the fact that an unrelated ancillary system should not affect the distance between the unital completely positive maps. Note however that Equation 3.1 does not make sense on the nose, because as it currently reads $d$ is a metric not only on the unital completely positive maps between $A$ and $B$, but also simultaneously on their amplifications. In [GLN05] and [BV25] it is expected that an analogue of $d$ exists on $\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ for every $n\in\mathbb{N}$. For our purposes this need not be the case, and we therefore define stability as a property of a sequence of metrics.
@@
Let $A$ and $B$ be unital $C^{*}$-algebras, and $\{d_{n}\}_{n\in\mathbb{N}}$ be sequence of extended metrics where each $d_{n}$ is an extended metric on
@@
$d_{n}(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)=d_{1}(F,G)$ (3.2)
@@
In Definition 4.13 we will introduce the set $\operatorname{TC}_{\tau}(A,B)$ of trace channels from $A$ to $B$, where $A$ is a unital $C^{*}$-algebra, and $B$ is a $C^{*}$-algebra with a faithful trace $\tau$. These form a particular subset of the completely positive maps from $A$ to $B$. We will then instead analogously define stability for a sequence $\{d_{n}\}_{n\in\mathbb{N}}$, where $d_{n}$ is an extended metric on
$\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B).$
The second property is chaining, which aims to capture the idea that if one divides a process into smaller steps, the total error should be smaller than the sum of the step-wise errors. Mathematically, it is defined as follows.
@@
Let $A$, $B$, and $C$ be unital $C^{*}$-algebras. We say that the three metrics $d_{1}$, $d_{2}$, and $d_{3}$ defined on $\operatorname{UCP}(A,B)$, $\operatorname{UCP}(B,C)$, and $\operatorname{UCP}(A,C)$ respectively, satisfy chaining if given $F_{1},F_{2}\in\operatorname{UCP}(A,B)$, and $G_{1},G_{2}\in\operatorname{UCP}(B,C)$, the following inequality is satisfied:
$d_{3}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq d_{2}(G_{1},G_{2})+d_{1}(F_{1},F_{2}).$ (3.3)
@@
Once again we will want to work with trace channels. It will however turn out that the composition of two trace channels is not necessarily a trace channel, hence we will in the sequel need to be careful when generalizing the definition of chaining to include trace channels as well. We will get back to this in Section 5.2.
In [GLN05] and [BV25] they discuss different ways of inducing metrics on $\operatorname{UCP}(A,B)$ given a metric on $S(A)$, or a metric on $S(A\otimes B^{\rm op})$ (recall that they only consider matrix algebras, hence there is a unique tensor product of $C^{*}$-algebras in this case). However, instead of starting with a metric on any of these state spaces, we start with noncommutative geometric data in the form of a seminorm. One way of inducing an extended metric on $\operatorname{UCP}(A,B)$ discussed in [GLN05, BV25] is as follows.
@@
$D_{L}(F,G):=\sup\limits_{\psi\in S(B)}\operatorname{mk}_{L}(F^{*}\psi,G^{*}\psi).$ (3.4)

### v2

@@ 2.2. Lipschitz seminorms and spectral triples
(2.1) $\operatorname{mk}_{L}(\phi,\psi):=\sup\{|\phi(a)-\psi(a)|:a\in A,L(a)\leq 1\}$
It is natural to wonder when $\operatorname{mk}_{L}$ is a genuine metric on $S(A)$. A result in this direction is that if $\ker(L)=\mathbb{C}1_{A}$ and
$\sup\big\{\inf\{\|a-\lambda 1_{A}\|:\lambda\in\mathbb{C}\}:a\in A,L(a)\leq 1\big\}<\infty,$
then $S(A)$ has finite diameter with respect to $\operatorname{mk}_{L}$ (cf. [Rie98, Proposition 1.6]). When $L$ is a Lipschitz seminorm on $A$, there has been a lot of interest in deciding when $\operatorname{mk}_{L}$ metrizes the weak-topology on $S(A)$, that is, when the topology induced from $\operatorname{mk}_{L}$ coincides with the weak-topology. When this happens, $(A,L)$ is called a compact quantum metric space, and it holds that $\ker(L)=\mathbb{C}1_{A}$ and $S(A)$ has finite diameter with respect to $\operatorname{mk}_{L}$.
@@
A unital spectral triple is even when the Hilbert space $H$ is equipped with a $\mathbb{Z}/2\mathbb{Z}$-grading operator $\gamma\colon H\to H$ such that $\pi(a)\colon H\to H$ is an even operator for all $a\in\mathcal{A}$ and $\partial\colon{\rm Dom}(\partial)\to H$ is odd. Otherwise we say the unital spectral triple is odd.
@@
(2.2) $L_{\partial}(a)=\begin{cases}\|[\partial,a]\|&\text{if $a\in\mathcal{A}$}\\ \infty&\text{if $a\in A\setminus\mathcal{A}$},\end{cases}$
@@
Constructing a Lipschitz seminorm from a spectral triple highlights an important subtlety: We set the domain ${\rm Dom}(L_{\partial})$ to be equal to $\mathcal{A}$, even if it is possible that the quantity $\|[\partial,a]\|<\infty$ for some $a\in A\setminus\mathcal{A}$. Despite this, we still set $L_{\partial}(a)=\infty$ in this case. All this is to say that the domain of $L_{\partial}$ need not agree with the maximal domain for the derivation $a\mapsto[\partial,a]$.
@@
(2.3) $\begin{split}(L_{A}\otimes 1_{B})(z)&:=\sup\{L_{A}((\operatorname{id}_{A}\otimes\psi)(z))\colon\psi\in S(B)\}\\ (1_{A}\otimes L_{B})(z)&:=\sup\{L_{B}((\phi\otimes\operatorname{id}_{B})(z))\colon\phi\in S(A)\}\\ L_{A\otimes B}&:=L_{A}\otimes 1_{B}+1_{A}\otimes L_{B}\end{split}$
with ${\rm Dom}(L_{A}\otimes 1_{B})={\rm Dom}(L_{A})\odot B$, ${\rm Dom}(1_{A}\otimes L_{B})=A\odot{\rm Dom}(L_{B})$, and ${\rm Dom}(L_{A}\otimes L_{B})={\rm Dom}(L_{A})\odot{\rm Dom}(L_{B})$, all these domains being dense in $A\otimes B$.
Assume now that $L_{A}$ and $L_{B}$ are Lipschitz seminorms. Using the fact that $L_{A}$ and $L_{B}$ are both *-invariant, it is not difficult to verify that the three seminorms just defined are *-invariant. Moreover, we easily verify that $\mathbb{C}\cdot(1_{A}\otimes 1_{B})$ is contained in the kernel of all three seminorms. Thus they are all Lipschitz seminorms. Note however that $\ker(L_{A}\otimes 1_{B})$ is generally larger than $\mathbb{C}\cdot(1_{A}\otimes 1_{B})$, as it at the very least contains $\mathbb{C}\cdot(1_{A}\otimes{\rm Dom}(L_{B}))$. The analogous statement holds for $1_{A}\otimes L_{B}$. Moreover, $L_{A\otimes B}(x)=L_{B\otimes A}(\Sigma(x))$ for all $x\in A\otimes B$.
@@
2.3. Reduced twisted group C*-algebras and length functions
Suppose $G$ is a discrete group and $\sigma\in Z^{2}(G,\mathbb{T})$ is a $2$-cocycle. The *-algebra $C_{c}(G,\sigma)$ consists of the complex functions on $G$ with finite support, equipped with convolution and involution defined by
$\delta_{g}(x)=\begin{cases}1&x=g,\\ 0&x\neq g.\end{cases}$
for $f\in C_{c}(G,\sigma)$, $\xi\in\ell^{2}(G)$ and $x\in G$. This defines a faithful *-representation of the opposite algebra $C^{*}_{r}(G,\sigma)^{\rm op}$.
Let $P(G)$ denote the positive definite functions on $G$, and set
(2.5) $P_{1}(G):=\{\phi\in P(G):\phi(1_{G})=1\}.$
is a completely positive map such that $\|M_{\phi}\|=\phi(1_{G})$ (resp. a unital completely positive map), see for instance [BO08, Theorem 2.5.11] (when $\sigma=1$) or [BC16, Lemma 2.6].
Definition 2.11 ([Con89]).
  1. (1)
    $l(s)=0$ if $s=1_{G}$,
  2. (2)
    $l(s^{-1})=l(s)$ for all $s\in G$, and
  3. (3)
    $l(st)\leq l(s)+l(t)$ for all $s,t\in G$.
Suppose moreover that $G$ is countable and the length function $l$ is proper, that is, for every $r\geq 0$, the set $l^{-1}([0,r])$ is a finite set. Such a length function gives rise to an essentially self-adjoint unbounded operator $\partial_{l}$ on $\ell^{2}(G)$ given by the linear extension of
$\begin{split}\partial_{l}\colon C_{c}(G)&\to\ell^{2}(G)\\ \delta_{s}&\mapsto l(s)\delta_{s},\text{ for }s\in G.\end{split}$
As $\partial_{l}$ is essentially self-adjoint, we may take its self-adjoint closure, which we will also denote by $\partial_{l}$ to ease notation. We thus obtain a spectral triple $(C^{*}_{r}(G,\sigma),\ell^{2}(G),\partial_{l})$, where $C^{*}_{r}(G,\sigma)$ is faithfully represented on $\ell^{2}(G)$ with the left regular representation of $C_{c}(G,\sigma)$ and the domain of $\partial_{l}$ is the image of $C_{c}(G,\sigma)$. As a result, we may define a seminorm $L_{\partial_{l}}$ through (2.2). We furthermore obtain a spectral triple $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$, where $C^{*}_{r}(G,\sigma)^{\rm op}$ is faithfully represented on $\ell^{2}(G)$ with the right regular representation of $C_{c}(G,\sigma)$. Note that a straightforward calculation such as in [Rie02, p. 614] shows that for any $f$ in $C_{c}(G,\sigma)$, the commutator $\partial_{l},\lambda^{\sigma}(f)]$ belongs to the reduced $\sigma$-twisted $C^{*}$-crossed product $\ell^{\infty}(G)\rtimes_{{\rm red},\sigma}G$, being here identified with the $C^{*}$-subalgebra of $B(\ell^{2}(G))$ generated by $\ell^{\infty}(G)$ (acting as multiplication operators on $\ell^{2}(G)$) and $\{\lambda^{\sigma}_{g}:g\in G\}$.
We will later need the following property of multipliers associated with $P(G)$ with respect to the Lipschitz seminorm $L_{\partial_{l}}$ on $C^{*}_{r}(G,\sigma)$.
Lemma 2.12.
Let $G$ be a countable discrete group, $l\colon G\to[0,\infty)$ be a proper length function, and $\sigma\in Z^{2}(G,\mathbb{T})$. Suppose $\phi\in P(G)$. Then $M_{\phi}$ extends to a completely positive multiplier $T^{\phi}$ of the twisted crossed product $\ell^{\infty}(G)\rtimes_{\mathrm{red},\sigma}G$, and we have
Proof.
The fact that $M_{\phi}$ extends to a completely positive map $T^{\phi}$ on $\ell^{\infty}(G)\rtimes_{\mathrm{red},\sigma}G$ satisfying
(2.6) $T^{\phi}(\psi\lambda^{\sigma}_{g})=\phi(g)\psi\lambda^{\sigma}_{g}\quad\text{ for all }\psi\in\ell^{\infty}(G)\text{ and }g\in G,$
Hence, for each $f\in C_{c}(G,\sigma)$ we get
$=\|T^{\phi}([\partial_{l},\lambda^{\sigma}(f)])\|$
$\leq\|T^{\phi}\|L_{\partial_{l}}(\lambda^{\sigma}(f))$
$=\|M_{\phi}\|\,L_{\partial_{l}}(\lambda^{\sigma}(f)).$
3. Discussions on inducing metrics on unital completely positive maps
Throughout this section, $A$ and $B$ will be unital $C^{*}$-algebras, and we will be considering extended metrics on the set $\operatorname{UCP}(A,B)$ of unital completely positive maps from $A$ to $B$.
@@ 3. Discussions on inducing metrics on unital completely positive maps
As mentioned in the introduction, the first property an extended metric on $\operatorname{UCP}(A,B)$ should satisfy is stability. In both [GLN05] and [BV25], a metric $d$ on $\operatorname{UCP}(A,B)$ is said to be stable if
(3.1) $d(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)=d(F,G)$
@@
holds for every $n\in\mathbb{N}$ and $F,G\in\operatorname{UCP}(A,B)$. Note that (3.1) does not make sense on the nose, because as it currently reads $d$ is a metric not only on the unital completely positive maps between $A$ and $B$, but also simultaneously on their amplifications. In [GLN05] and [BV25] it is expected that an analogue of $d$ exists on $\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ for every $n\in\mathbb{N}$. For our purposes this need not be the case, and we therefore define stability as a property of a sequence of extended metrics.
@@
Let $A$ and $B$ be unital $C^{*}$-algebras, and $\{d_{n}\}_{n\in\mathbb{N}}$ be a sequence of extended metrics where for each $n\geq 1$, $d_{n}$ is an extended metric on
@@
(3.2) $d_{n}(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)=d_{1}(F,G)$
@@
We next formalize the property of chaining from [GLN05] mathematically as follows.
@@
Let $A$, $B$ and $C$ be unital $C^{*}$-algebras. We say that extended metrics $d_{1}$, $d_{2}$ and $d_{3}$, defined on $\operatorname{UCP}(A,B)$, $\operatorname{UCP}(B,C)$ and $\operatorname{UCP}(A,C)$, respectively, satisfy chaining if given any $F_{1},F_{2}\in\operatorname{UCP}(A,B)$, and any $G_{1},G_{2}\in\operatorname{UCP}(B,C)$, the following inequality is satisfied:
(3.3) $d_{3}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq d_{2}(G_{1},G_{2})+d_{1}(F_{1},F_{2}).$
@@
Note that in a lot of circumstances it might be too strict to require that (3.3) holds for all unital completely positive maps. Indeed, even in the matrix algebra case, some restrictions are imposed on at least one of the maps involved in the analogue of (3.3), see for instance [GLN05, p. 5] and [BLZ23, p. 4].
In [GLN05] and [BV25], methods for inducing metrics on quantum channels between matrix algebras $A$ and $B$ from metrics on $S(A)$ (or on $S(A\otimes B)$) are discussed, using the canonical identification of these state spaces with the corresponding spaces of density matrices. However, instead of starting with a metric on any of these state spaces, we will start with noncommutative geometric data in the form of a seminorm.
First, if $L$ is a seminorm on $A$, we can use the associated Monge-Kantorovič extended metric $\operatorname{mk}_{L}$ on $S(A)$, cf. (2.1), to induce an extended metric $d_{L}$ on $\operatorname{UCP}(A,B)$ as follows.
@@
(3.4) $d_{L}(F,G):=\sup\limits_{\psi\in S(B)}\operatorname{mk}_{L}(F^{*}\psi,G^{*}\psi)=\sup\limits_{\psi\in S(B)}\operatorname{mk}_{L}(\psi\circ F,\psi\circ G)$
for all $F,G\in\operatorname{UCP}(A,B)$. An extended metric induced in this manner will be called a pullback-induced extended metric from a seminorm.